# Learning a distance measure from the information-estimation geometry of data

### v1

@@ Abstract
We introduce the Information-Estimation Metric (IEM), a novel form of distance function derived from an underlying continuous probability density over a domain of signals. The IEM is rooted in a fundamental relationship between information theory and estimation theory, which links the log-probability of a signal with the errors of an optimal denoiser, applied to noisy observations of the signal. In particular, the IEM between a pair of signals is obtained by comparing their denoising error vectors over a range of noise amplitudes. Geometrically, this amounts to comparing the score vector fields of the blurred density around the signals over a range of blur levels. We prove that the IEM is a valid global metric and derive a closed-form expression for its local second-order approximation, which yields a Riemannian metric. For Gaussian-distributed signals, the IEM coincides with the Mahalanobis distance. But for more complex distributions, it adapts, both locally and globally, to the geometry of the distribution. In practice, the IEM can be computed using a learned denoiser (analogous to generative diffusion models) and solving a one-dimensional integral. To demonstrate the value of our framework, we learn an IEM on the ImageNet database. Experiments show that this IEM is competitive with or outperforms state-of-the-art supervised image quality metrics in predicting human perceptual judgments.
@@ 1 Introduction
A natural opportunity for developing such a metric arises from the concept of coding efficiency. Biological sensory systems are believed to decompose incoming signals in a manner that maximizes the transmission of information about those signals, subject to biological constraints (e.g., noise, metabolic cost) (Attneave, 1954; Barlow, 1961; Laughlin, 1981; van Hateren, 1992; Atick & Redlich, 1992; Olshausen & Field, 1996; Barlow, 2001; Simoncelli & Olshausen, 2001). Put differently, sensory pathways function as communication channels optimized for natural signals, implying that our ability to discriminate between natural signals depends on their statistical properties. Indeed, for one-dimensional sensory attributes, previous work has shown that perceptual sensitivity to small signal perturbations increases with the probability of the signal (Laughlin, 1981; Ganguli & Simoncelli, 2014; Wei & Stocker, 2017). However, in the multivariate setting, such as color discrimination (i.e., detecting changes in hue or saturation), humans exhibit complex patterns of sensitivity that vary with the direction of the signal’s perturbation (MacAdam, 1942). This leaves us with a conundrum: how can a probability density, which is a scalar function, induce a Riemannian metric, let alone a global distance function between any pair of signals in the domain?
@@ 2 The Information-Estimation Metric
Venkat & Weissman (2012) proved that $\log{p_{{\mathbf{y}}_{{\Gamma}}}({\mathbf{y}}_{\Gamma})}$ for any fixed $\smash{{\Gamma}>0}$ can be expressed in terms of the denoising error vectors of the minimum mean square error (MMSE) estimator of $\mathbf{x}$} from ${\mathbf{y}}_{\gamma}$, $\mathbb{E}\mathopen{}\mathclose{{\left[{\mathbf{x}}\,|\,{\mathbf{y}}_{\gamma}}}\right]$, integrated across all SNR levels $\gamma\in[0,{\Gamma}]$. Formally,
@@
where this equality holds with probability one (i.e., almost surely). Importantly, this relationship holds pointwise for any realization of the signal ${\mathbf{x}}={\bm{x}}$ and the noise ${\mathbf{w}}_{\gamma}={\bm{w}}_{\gamma}$. Equation 2, which we refer to as the pointwise I-MMSE formula, is a generalization of the I-MMSE formula (Guo et al., 2005), whose roots date back to de Bruijn’s identity from the 1950s (Stam, 1959, see Sec. A.2 for more detailed background). When $\Gamma$$\to\inftyEq$. 2 expresses the log-density of the original signal, $\log p_{\mathbf{x}}({\mathbf{x}})$, in terms of the denoising errors at all $\gamma\in[0,\infty)$.
@@
where the gradient on the right-hand side is taken with respect to ${\mathbf{y}}_{\gamma}$. Substituting this formula into Eq. 2, we now see that $\log p_{{\mathbf{x}}}({\mathbf{x}})$ (a scalar) can be decomposed in terms of the local geometry of $\log{p_{{\mathbf{y}}_{\gamma}}}({\mathbf{y}}_{\gamma})$, particularly the gradients $\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})$, at all SNR levels $\gamma$. We refer to this decomposition as the information-estimation geometry of the density $p_{{\mathbf{x}}}$.
@@
${}-\frac{1}{2}\int_{0}^{{\Gamma}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\lVert\nabla\log{p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}_{1}+{\mathbf{w}}_{\gamma})}-{\mathbf{w}}_{\gamma}/\gamma}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lVert\nabla\log{p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}_{2}+{\mathbf{w}}_{\gamma})}-{\mathbf{w}}_{\gamma}/\gamma}}\right\rVert^{2}}}\right)\mathrm{d}\gamma.$ (4)
@@ 2.1 Local geometry
To gain insight into the properties of the IEM, we study its local behavior, namely the distance between a given signal $\bm{x}$} and its perturbation $\smash{{\bm{x}}+\bm{\epsilon}}$ for small $\bm{\epsilon}$. As for any distance, $\smash{\bm{\epsilon}=0}$ is a global minimum, and we can express the quadratic expansion of the IEM in $\bm{\epsilon}$ as $\smash{\text{IEM}^{2}({\bm{x}},{\bm{x}}+\bm{\epsilon},{\Gamma})=\bm{\epsilon}^{\top}{\bm{G}}({\bm{x}},{\Gamma})\bm{\epsilon}+o(\mathopen{}\mathclose{{\left\lVert\bm{\epsilon}}}\right\rVert^{2})}$. The positive-definite matrix ${\bm{G}}({\bm{x}},{\Gamma})$ then acts as a local metric (in the Riemannian sense), but note that its relationship with the IEM is one-way: the IEM is not equivalent to the geodesic distance that corresponds to ${\bm{G}}({\mathbf{x}},{\Gamma})$. The local metric ${\bm{G}}({\bm{x}},{\Gamma})$ is characterized in the following theorem (see proof in Sec. C.2):
@@
The local Riemannian metric derived from the second-order Taylor expansion of the IEM is given by
@@ 2.2 Illustrative examples
First, consider a two-dimensional, two-mode Gaussian mixture model. To compute the IEM and the local metric ${\bm{G}}({\mathbf{x}},{\Gamma})$, we numerically solve the integrals in Defs. 1 and 5, using closed-form expressions for $\log{p_{{\mathbf{y}}_{\gamma}}}$ and related quantities (see details in Sec. E.4). To illustrate the global behavior of the IEM, we choose a reference point ${\bm{x}}_{\text{ref.}}$, and evaluate its distance from each of a uniform grid of points $\bm{x}$}. We then plot the resulting equidistant contours (Fig. 2, top row) , and compare with a unimodal Gaussian density to illustrate how the IEM adapts to the prior. The IEM clearly adapts to the density’s global geometry: the equidistant contours resemble the shape of the log-density contours. Interestingly, the regions delimited by equidistant contours can be disconnected: points belonging to one mode are closer to points belonging to the other mode than they are to points lying in between the modes. This is because the local curvature of the log-density can be similar in the vicinity of two points, even if their Euclidean distance is large.
Furthermore, we eigendecompose ${\bm{G}}({\bm{x}},{\Gamma})^{-\frac{1}{2}}$ at each point $\bm{x}$} on the grid, and draw an ellipse centered at $\bm{x}$}, whose axes and radii are the resulting eigenvectors and the corresponding eigenvalues, respectively. These ellipses represent the discrimination thresholds of the metric across space, which are inversely proportional to its local sensitivities. In other words, the ellipses illustrate the directions that require larger perturbations to induce the same change in distance. Figure 2 shows that the discrimination thresholds align with the direction of the local covariance, i.e., the metric ${\bm{G}}({\bm{x}},{\Gamma})$ behaves like a locally adaptive Mahalanobis metric. We also note that ${\bm{G}}({\bm{x}},{\Gamma})$ is more sensitive in the local minima of probability in between the modes, as illustrated by the ellipses with smaller radii (smaller discrimination thresholds). This is consistent with Thm. 2, as the signals lying between modes incur large denoising errors due to uncertainty about the mode they belong to.
@@
To further illustrate the influence of the density’s curvature on the IEM, we now consider the case where $p_{\mathbf{x}}$ is a two-dimensional Laplace distribution, formed by taking the product of one-dimensional Laplace densities. As shown in Fig. 2, this prior density induces discrimination thresholds that increase as they move away from the high-probability ridges that lie along the axes. From the point of view of Thm. 2, this reflects the fact that the Hessian matrices $\smash{\nabla^{2}\log{p_{{\mathbf{y}}_{\gamma}}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}$ (specifically, their negative eigenvalues) decrease in magnitude away from the axes. For this sparse and heavy-tailed distribution, curvature is correlated with probability, so that discrimination thresholds are larger in low-probability regions, consistent with predictions from prior work on efficient coding (Ganguli & Simoncelli, 2014). From the global behavior of the distance, we also see that the equidistant contour lines tend to cluster around the axes: under a sparse prior such as the Laplace density, flipping the sign of one or several coordinates of $\bm{x}$} (landing on the other side of the high-probability ridge) incurs a large cost as measured by the IEM.
@@ 2.3 Generalized Information-Estimation Metric
For $f({\mathbf{z}}_{\gamma},\gamma)={\mathbf{z}}_{\gamma}$, $\text{IEM}_{f}$ recovers the IEM from Def. 1. Moreover, when $\smash{f({\mathbf{z}}_{\gamma},\gamma)}$ satisfies $\smash{f^{\prime}({\mathbf{z}}_{\gamma},\gamma)=0}$ if and only if $\smash{{\mathbf{z}}_{\gamma}=0}$, we have $\smash{\text{IEM}_{f}({\bm{x}}_{1},{\bm{x}}_{2})=0}$ if and only if $\smash{{\bm{x}}_{1}={\bm{x}}_{2}}$ (positive definiteness). Unlike the IEM, however, the $\text{IEM}_{f}$ is generally not a proper metric, as it may violate the symmetry or the triangle inequality axioms. This may or may not be considered a limitation, depending on the intended application of the distance. Definition 2 can be extended to any non-anticipative functional $\smash{f(\{{\mathbf{z}}_{\gamma^{\prime}}\}_{\gamma^{\prime}=0}^{\gamma},\gamma)}$ whose input is the entire history of the process ${\mathbf{z}}_{\gamma^{\prime}}$ up to SNR $\gamma$. In this case, $f^{\prime}$ is a Dupire derivative (Dupire, 2009; Cont & Fournie, 2010).
@@ 3.2 Predicting mean opinion scores
We consider several standard full-reference image quality assessment benchmarks, including TID2013 (Ponomarenko et al., 2015), CSIQ (Larson & Chandler, 2010), and LIVE (Sheikh et al., 2006). Since the learned denoiser model is suited for images of size $256\times 256$, we adjust the resolution of the images in the considered databases by first center-cropping each image to the length of its shorter edge, and then resizing it to $256\times 256$. We compare against PSNR, SSIM (Wang et al., 2004), VIF (Sheikh & Bovik, 2006), MAD (Larson & Chandler, 2010), FSIM (Zhang et al., 2011), GMSD (Xue et al., 2014), NLPD (Laparra et al., 2016), PieAPP (Prashnani et al., 2018), LPIPS (Zhang et al., 2018), DISTS (Ding et al., 2022), and TOPIQ (Chen et al., 2024). We find that the IEM with $\Gamma=1/4$ yields surprisingly strong results, even though it is computed solely based on denoising errors and is not exposed to human labels. Indeed, as shown in Fig. 4, this same choice of $\Gamma$ produces a strikingly high Spearman’s rank correlation coefficient (SRCC) with the human MOS across all of the aforementioned datasets. Additional performance measures demonstrate similar trends, so we report them in Figs. 7 and 6 in Sec. D.2.
To illustrate the differences between the IEM and the compared distance measures, we present in Fig. 3 several example images for which the IEM rankings differ the most from those of the compared methods. For each reference image in the dataset, we rank all of its distorted counterparts according to each distance measure. We then compute the difference between the ranks assigned by the IEM and those assigned by each compared method. From these differences, we take the maximum and minimum values, and sum their absolute magnitudes to quantify the degree of disagreement. Finally, we display the reference and distorted images that achieve the largest disagreement. This systematic procedure for comparing image similarity models on a given dataset is analogous to the group maximum differentiation competition of Ma et al. (2018). The results show that VIF, FSIM, PieAPP, and LPIPS can assign smaller distances to image pairs that are perceptually distinguishable, whereas the IEM correctly detects that the images are different. In comparison, MAD and DISTS tend to disagree with the IEM in cases involving perceptually noticeable distortions. For example, DISTS appears to favor noise over blur, whereas the IEM shows the opposite preference.
@@
We further evaluate our distance measures on the TQD textures dataset (Ding et al., 2022). Unlike the previously considered benchmarks, which contain general natural images, this dataset consists of texture images (e.g., leaves or brick walls), paired both with visually similar textures and with distorted versions of the same texture (e.g., Gaussian blur, JPEG compression). In this setting, human observers are expected to judge two images of the same texture as more similar than a clean-distorted pair. Thus, the TQD benchmark assesses whether a perceptual distance measure produces scores consistent with human perception even when the compared images are substantially different in terms of their Euclidean distance. As in the previous experiments, we use the $256\times 256$ denoiser model with the same preprocessing to resize the images.

### v2

@@ Abstract
We introduce the Information-Estimation Metric (IEM), a novel form of distance function derived from an underlying continuous probability density over a domain of signals. The IEM is rooted in a fundamental relationship between information theory and estimation theory, which links the log-probability of a signal with the errors of an optimal denoiser, applied to noisy observations of the signal. In particular, the IEM between a pair of signals is obtained by comparing their denoising error vectors over a range of noise amplitudes. Geometrically, this amounts to comparing the score vector fields of the blurred density around the signals over a range of blur levels. We prove that the IEM is a valid global distance metric and derive a closed-form expression for its local second-order approximation, which yields a Riemannian metric. For Gaussian-distributed signals, the IEM coincides with the Mahalanobis distance. But for more complex distributions, it adapts, both locally and globally, to the geometry of the distribution. In practice, the IEM can be computed using a learned denoiser (analogous to generative diffusion models) and solving a one-dimensional integral. To demonstrate the value of our framework, we learn an IEM on the ImageNet database. Experiments show that this IEM is competitive with or outperforms state-of-the-art supervised image quality metrics in predicting human perceptual judgments.
@@ 1 Introduction
A natural opportunity for developing such a metric arises from the concept of coding efficiency. Biological sensory systems are believed to decompose incoming signals in a manner that maximizes the transmission of information about those signals, subject to biological constraints (e.g., noise, metabolic cost) (Attneave, 1954; Barlow, 1961; Laughlin, 1981; van Hateren, 1992; Atick and Redlich, 1992; Olshausen and Field, 1996; Barlow, 2001; Simoncelli and Olshausen, 2001). Put differently, sensory pathways function as communication channels optimized for natural signals, implying that our ability to discriminate between natural signals depends on their statistical properties. Indeed, for one-dimensional sensory attributes, previous work has shown that perceptual sensitivity to small signal perturbations increases with the probability of the signal (Laughlin, 1981; Ganguli and Simoncelli, 2014; Wei and Stocker, 2017). However, in the multivariate setting, such as color discrimination (i.e., detecting changes in hue or saturation), humans exhibit complex patterns of sensitivity that vary with the direction of the signal’s perturbation (MacAdam, 1942). This leaves us with a conundrum: how can a probability density, which is a scalar function, induce a Riemannian metric, let alone a global distance function between any pair of signals in the domain?
@@ 2 The Information-Estimation Metric
Venkat and Weissman (2012) proved that $\log{p_{{\mathbf{y}}_{{\Gamma}}}({\mathbf{y}}_{\Gamma})}$ for any fixed $\smash{{\Gamma}>0}$ can be expressed in terms of the denoising error vectors of the minimum mean square error (MMSE) estimator of $\mathbf{x}$} from ${\mathbf{y}}_{\gamma}$, $\mathbb{E}\mathopen{}\mathclose{{\left[{\mathbf{x}}\,|\,{\mathbf{y}}_{\gamma}}}\right]$, integrated across all SNR levels $\gamma\in[0,{\Gamma}]$. Formally,
@@
where this equality holds with probability one (almost surely), i.e., it holds pointwise for almost every realization of the signal $\smash{{\mathbf{x}}={\bm{x}}}$ and of the Wiener process trajectory $\smash{\{{\mathbf{w}}_{\gamma}={\bm{w}}_{\gamma}\}_{\gamma=0}^{{\Gamma}}}$. Equation 2, which we refer to as the pointwise I-MMSE formula, is a generalization of the I-MMSE formula (Guo et al., 2005), whose roots date back to de Bruijn’s identity from the 1950s (Stam, 1959, see Sec. A.2 for more detailed background). When $\Gamma$$\to\inftyEq$. 2 expresses the log-density of the original signal, $\log p_{\mathbf{x}}({\mathbf{x}})$, in terms of the denoising errors at all $\gamma\in[0,\infty)$.
@@
where the gradient on the right-hand side is taken w.r.t. ${\mathbf{y}}_{\gamma}$. Substituting this formula into Eq. 2, we now see that $\log p_{{\mathbf{x}}}({\mathbf{x}})$ (a scalar) can be decomposed in terms of the local geometry of $\log{p_{{\mathbf{y}}_{\gamma}}}({\mathbf{y}}_{\gamma})$, particularly the gradients $\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})$, at all SNR levels $\gamma$. We refer to this decomposition as the information-estimation geometry of the density $p_{{\mathbf{x}}}$.
@@
${}-\frac{1}{2}\int_{0}^{{\Gamma}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\lVert\nabla\log{p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}_{1}+{\mathbf{w}}_{\gamma})}+{\mathbf{w}}_{\gamma}/\gamma}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lVert\nabla\log{p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}_{2}+{\mathbf{w}}_{\gamma})}+{\mathbf{w}}_{\gamma}/\gamma}}\right\rVert^{2}}}\right)\mathrm{d}\gamma.$ (4)
@@ 2.1 Local geometry
To gain insight into the properties of the IEM, we study its local behavior, namely the distance between a given signal $\bm{x}$} and its perturbation $\smash{{\bm{x}}+\bm{\epsilon}}$ for small $\bm{\epsilon}$. As for any distance, $\smash{\bm{\epsilon}=0}$ is a global minimum, and we can express the quadratic expansion of the IEM in $\bm{\epsilon}$ as $\smash{\text{IEM}^{2}({\bm{x}},{\bm{x}}+\bm{\epsilon},{\Gamma})=\bm{\epsilon}^{\top}{\bm{G}}({\bm{x}},{\Gamma})\bm{\epsilon}+o(\mathopen{}\mathclose{{\left\lVert\bm{\epsilon}}}\right\rVert^{2})}$. The positive-definite matrix ${\bm{G}}({\bm{x}},{\Gamma})$ then acts as a local metric (in the Riemannian sense), but note that its relationship with the IEM is one-way: the IEM is not equivalent to the geodesic distance that corresponds to ${\bm{G}}({\bm{x}},{\Gamma})$. The local metric ${\bm{G}}({\bm{x}},{\Gamma})$ is characterized in the following theorem (see proof in Sec. C.2):
@@
The local Riemannian metric derived from the second-order Taylor expansion of the squared IEM is given by
@@ 2.2 Illustrative examples
First, consider a two-dimensional, two-mode Gaussian mixture model. To compute the IEM and the local metric ${\bm{G}}({\bm{x}},{\Gamma})$, we numerically solve the integrals in Defs. 1 and 5, using closed-form expressions for $\log{p_{{\mathbf{y}}_{\gamma}}}$ and related quantities (see details in Sec. E.4). To illustrate the global behavior of the IEM, we choose a reference point ${\bm{x}}_{\text{ref.}}$, and evaluate its distance from each of a uniform grid of points $\bm{x}$}. We then plot the resulting equidistant contours (Fig. 2, top row) , and compare with a unimodal Gaussian density to illustrate how the IEM adapts to the prior. The IEM clearly adapts to the density’s global geometry: the equidistant contours resemble the shape of the log-density contours. Interestingly, the regions delimited by equidistant contours can be disconnected: points belonging to one mode are closer to points belonging to the other mode than they are to points lying in between the modes. This is because the local curvature of the log-density can be similar in the vicinity of two points, even if their Euclidean distance is large.
Furthermore, we eigendecompose ${\bm{G}}({\bm{x}},{\Gamma})^{-\frac{1}{2}}$ at each point $\bm{x}$} on the grid, and draw an ellipse centered at $\bm{x}$}, whose axes and radii are the resulting eigenvectors and the corresponding eigenvalues, respectively. These ellipses represent the discrimination thresholds of the metric across space, which are inversely proportional to its local sensitivities. In other words, the ellipses illustrate the directions that require larger perturbations to induce the same change in distance. Figure 2 shows that the discrimination thresholds align with the direction of the local covariance, i.e., the metric ${\bm{G}}({\bm{x}},{\Gamma})$ behaves like a locally adaptive Mahalanobis metric. We also note that ${\bm{G}}({\bm{x}},{\Gamma})$ is more sensitive in the local minima of probability in between the modes, as illustrated by the ellipses with smaller radii (smaller discrimination thresholds). This is consistent with Thm. 2, as the signals lying between modes incur large denoising errors due to the uncertainty about the mode they belong to.
@@
To further illustrate the influence of the density’s curvature on the IEM, we now consider the case where $p_{\mathbf{x}}$ is a two-dimensional Laplace distribution, formed by taking the product of one-dimensional Laplace densities. As shown in Fig. 2, this prior density induces discrimination thresholds that increase as they move away from the high-probability ridges that lie along the axes. From the point of view of Thm. 2, this reflects the fact that the Hessian matrices $\smash{\nabla^{2}\log{p_{{\mathbf{y}}_{\gamma}}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}$ (specifically, their negative eigenvalues) decrease in magnitude away from the axes. For this sparse and heavy-tailed distribution, curvature is correlated with probability, so that discrimination thresholds are larger in low-probability regions, consistent with predictions from prior work on efficient coding (Ganguli and Simoncelli, 2014). From the global behavior of the distance, we also see that the equidistant contour lines tend to cluster around the axes: under a sparse prior such as the Laplace density, flipping the sign of one or several coordinates of $\bm{x}$} (landing on the other side of the high-probability ridge) incurs a large cost as measured by the IEM.
@@ 2.3 Generalized Information-Estimation Metric
For $f({\mathbf{z}}_{\gamma},\gamma)={\mathbf{z}}_{\gamma}$, $\text{IEM}_{f}$ recovers the IEM from Def. 1. Moreover, when $\smash{f({\mathbf{z}}_{\gamma},\gamma)}$ satisfies $\smash{f^{\prime}({\mathbf{z}}_{\gamma},\gamma)=0}$ if and only if $\smash{{\mathbf{z}}_{\gamma}=0}$, we have $\smash{\text{IEM}_{f}({\bm{x}}_{1},{\bm{x}}_{2})=0}$ if and only if $\smash{{\bm{x}}_{1}={\bm{x}}_{2}}$ (positive definiteness). Unlike the IEM, however, the $\text{IEM}_{f}$ is generally not a proper metric, as it may violate the symmetry or the triangle inequality axioms. This may or may not be considered a limitation, depending on the intended application of the distance. Definition 2 can be extended to any non-anticipative functional $\smash{f(\{{\mathbf{z}}_{\gamma^{\prime}}\}_{\gamma^{\prime}=0}^{\gamma},\gamma)}$ whose input is the entire history of the process ${\mathbf{z}}_{\gamma^{\prime}}$ up to SNR $\gamma$. In this case, $f^{\prime}$ is a Dupire derivative (Dupire, 2009; Cont and Fournie, 2010).
@@ 3.2 Predicting mean opinion scores
We consider several standard full-reference image quality assessment benchmarks, including TID2013 (Ponomarenko et al., 2015), CSIQ (Larson and Chandler, 2010), and LIVE (Sheikh et al., 2006). Since the learned denoiser model is suited for images of size $256\times 256$, we adjust the resolution of the images in the considered databases by first center-cropping each image to the length of its shorter edge, and then resizing it to $256\times 256$. We compare against PSNR, SSIM (Wang et al., 2004), VIF (Sheikh and Bovik, 2006), MAD (Larson and Chandler, 2010), FSIM (Zhang et al., 2011), GMSD (Xue et al., 2014), NLPD (Laparra et al., 2016), PieAPP (Prashnani et al., 2018), LPIPS (Zhang et al., 2018), DISTS (Ding et al., 2022), and TOPIQ (Chen et al., 2024). We find that the IEM with $\Gamma=1/4$ yields surprisingly strong results, even though it is computed solely based on denoising errors and is not exposed to human labels. Indeed, as shown in Fig. 4, this same choice of $\Gamma$ produces a strikingly high Spearman’s rank correlation coefficient (SRCC) with the human MOS across all of the aforementioned datasets. Additional performance measures demonstrate similar trends, so we report them in Figs. 8 and 7 in Sec. D.2.
To illustrate the differences between the IEM and the compared distance measures, we present in Fig. 3 several example images for which the IEM rankings differ the most from those of the compared methods. For each reference image in the dataset, we rank all of its distorted counterparts according to each distance measure. We then compute the difference between the ranks assigned by the IEM and those assigned by each compared method. From these differences, we take the maximum and minimum values, and sum their absolute magnitudes to quantify the degree of disagreement. Finally, we display the reference and distorted images that achieve the largest disagreement. This systematic procedure for comparing image similarity models on a given dataset is analogous to the maximum differentiation competition (Wang and Simoncelli, 2008). The results show that VIF, FSIM, PieAPP, and LPIPS can assign smaller distances to image pairs that are perceptually distinguishable, whereas the IEM correctly detects that the images are different. In comparison, MAD and DISTS tend to disagree with the IEM in cases involving perceptually noticeable distortions. For example, DISTS appears to favor noise over blur, whereas the IEM shows the opposite preference.
@@
We further evaluate our distance measures on the TQD textures dataset (Ding et al., 2022). Unlike the previously considered benchmarks, which contain general natural images, this dataset consists of texture images (e.g., leaves or brick walls), paired both with visually similar textures and with distorted versions of the same texture (e.g., Gaussian blur, JPEG compression). In this setting, human observers are expected to judge two images of the same texture as more similar to each other than a clean and distorted pair. Thus, the TQD benchmark assesses whether a perceptual distance measure produces scores consistent with human perception even when the compared images are substantially different in terms of their Euclidean distance. As in the previous experiments, we use the $256\times 256$ denoiser model with the same preprocessing to resize the images.
@@
3.3 Maximum differentiation competition against the PSNR measure
Figure 5: Visual illustration of the maximum differentiation competition. We corrupt a given reference image with random noise to produce a distorted image with $\text{PSNR}=10\text{dB}$. This distorted image lies on the surface of a hypersphere in $\mathbb{R}^{d}$ centered at the reference image, with radius equal to the Euclidean norm of the added noise, as illustrated above. Starting from the distorted image, we then minimize or maximize the perceptual distance to the reference image while constraining the PSNR to $10\text{dB}$. Even under such a restrictive constraint, minimizing the IEM yields artifact-free images that are perceptually similar to the reference image, whereas minimizing state-of-the-art supervised perceptual metrics such as DISTS yields unrealistic images with noticeable artifacts. Furthermore, the results obtained by maximizing the IEM support our theoretical analysis in Sec. 2, demonstrating that the IEM is most sensitive to unstructured distortions (e.g., additive noise) that perturb the reference image outside the “data support.”
To further demonstrate the behavior of the IEM, we conduct a maximum differentiation competition against PSNR. Specifically, we minimize or maximize each perceptual distance (IEM, DISTS, etc.) via projected gradient descent or ascent, respectively, where the projection step during optimization constrains the PSNR of the distorted image to a fixed, pre-determined level. Figure 5 illustrates this experiment and compares the optimized images produced by the IEM and DISTS, using a PSNR constraint of $10\text{dB}$. Comparisons with additional perceptual distance measures on varying levels of PSNR constraints are shown in Figs. 9, 10, 11 and 12. Implementation details are disclosed in Sec. D.3.
Interestingly, we find that minimizing the IEM consistently yields high perceptual quality images that preserve the overall geometric structure of the reference image, even under very low PSNR constraints (see Figs. 5, 9, 10, 11 and 12). In contrast, minimizing other perceptual metrics produces images with unnatural artifacts, even for relatively high PSNR constraints (e.g., $25\text{dB}$). These results suggest that, unlike previous perceptual metrics, the IEM may serve as a stand-alone robust optimization objective (e.g., for solving inverse problems), and we encourage future work to explore this potential. Furthermore, maximizing the IEM reveals that it is most sensitive to unstructured noise perturbations that push an image off the “data support” (decreasing the image’s probability), consistent with our theoretical analysis in Sec. 2.